# How Data Mixing Shapes In-Context Learning: Asymptotic Equivalence for Transformers with MLPs

### v1

@@ Appendix D Conditional Gaussian equivalence
where $\bm{v}$} is the spiked direction from Lemma 4.11, and $\bm{\gamma}_{s,1},\dots,\bm{\gamma}_{s,r_{s}}$ are the spiked directions of $\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))$ as specified in Assumption 4.4. Let $\bm{P}$} and $\bm{P}$$_{\perp}$ denote the orthogonal projection matrices onto $\text{span}({\bm{S}})$ and its orthogonal complement, respectively. Then,
@@
${\bm{\Phi}}(s,\bm{\kappa}_{s})$ $:=\text{Cov}\left(\sigma(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{{\bm{Z}}}))\,\middle|\,s,\bm{\kappa}_{s}\right)-\bm{\Psi}(s,\bm{\kappa}_{s})\bm{\Psi}(s,\bm{\kappa}_{s})^{T}.$
@@
where the supremum is taken over $\tilde{{\bm{w}}}\in\{{\bm{w}}\in\mathbb{R}^{k}\mid\|{\bm{w}}\|=\mathcal{O}(1),\|{\bm{w}}\|_{\infty}=\mathcal{O}(k^{-\epsilon})\}$ for some $\epsilon>0$, and $\tilde{\bm{\xi}}\in\mathbb{R}^{d}$ satisfies $\|\tilde{\bm{\xi}}\|=1/\|\text{Cov}({\bm{x}}\mid s)\|^{1/2}$. Here, $k,d\to\infty$ such that $k/d^{2}\in\mathbb{R}$, as specified in Assumption 4.1.
@@
  • $\|\bm{\Delta}\|\to 0$ asymptotically (see Appendix C),
  • $\bm{P}$$_{\perp}$ is an orthogonal projection ($\|{\bm{P}}_{\perp}\|=1$),
@@
  • Assumption 4.4 ensures $\bm{P}$$_{\perp}$ eliminates spiked directions, leading to ${\bm{r}}\to\mathcal{N}(0,{\bm{I}}_{k})$ almost surely.
@@ Appendix E Equivalent polynomial model
$\left\|\mathbb{E}\left[\sigma(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{\bm{Z}}))\;\middle|\;s,\bm{\kappa}_{s}\right]-\mathbb{E}\left[\hat{\sigma}_{p}(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{\bm{Z}}))\;\middle|\;s,\bm{\kappa}_{s}\right]\right\|=o(1),$ (S34)
@@
$\left\|\text{Cov}\left(\sigma(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{\bm{Z}}))\;\middle|\;s,\bm{\kappa}_{s}\right)-\text{Cov}\left(\hat{\sigma}_{p}(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{\bm{Z}}))\;\middle|\;s,\bm{\kappa}_{s}\right)\right\|=o(1),$ (S36)
@@
Finally, invoking the bounds derived in Appendix C, along with Assumptions 4.4 and 4.5, we conclude that each term vanishes at the desired rate. These assumptions ensure that the step size $\eta$ and the influence of spiked directions in $\mathrm{Cov}(\text{vec}({\bm{H}}_{\bm{Z}}))$ remain controlled, which completes the proof.

### v2

@@ Appendix D Conditional Gaussian equivalence
where $\bm{v}$} is the spiked direction from Lemma 4.11, and $\bm{\gamma}_{s,1},\dots,\bm{\gamma}_{s,r_{s}}$ are the spiked directions of $\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))$ as specified in Assumption 4.4. Let $\bm{P}$} and $\bm{P}$$_{\perp}$ denote the orthogonal projection matrices onto $\text{span}({\bm{S}})$ and its orthogonal complement, respectively. Then,
@@
${\bm{\Phi}}(s,\bm{\kappa}_{s})$ $:=\mathrm{Cov}\left(\sigma(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{{\bm{Z}}}))\,\middle|\,s,\bm{\kappa}_{s}\right)-\bm{\Psi}(s,\bm{\kappa}_{s})\bm{\Psi}(s,\bm{\kappa}_{s})^{T}.$
@@
where the supremum is taken over $\tilde{{\bm{w}}}\in\{{\bm{w}}\in\mathbb{R}^{k}\mid\|{\bm{w}}\|_{2}=\mathcal{O}(1),\|{\bm{w}}\|_{\infty}=\mathcal{O}(k^{-\epsilon})\}$ for some $\epsilon>0$, and $\tilde{\bm{\xi}}\in\mathbb{R}^{d}$ satisfies $\|\tilde{\bm{\xi}}\|_{2}=1/\|\mathrm{Cov}({\bm{x}}\mid s)\|_{2}^{1/2}$. Here, $k,d\to\infty$ such that $k/d^{2}\in\mathbb{R}$, as specified in Assumption 4.1.
@@
  • $\eta\|\bm{\Delta}\|_{2}\to 0$ asymptotically (see Appendix C),
  • $\bm{P}$$_{\perp}$ is an orthogonal projection ($\|{\bm{P}}_{\perp}\|_{2}=1$),
@@
  • Assumption 4.4 ensures $\bm{P}$$_{\perp}$ eliminates spiked directions, leading to ${\bm{r}}\to\mathcal{N}(0,{\bm{I}}_{k})$.
@@ Appendix E Equivalent polynomial model
$\left\|\mathbb{E}\left[\sigma(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{\bm{Z}}))\;\middle|\;s,\bm{\kappa}_{s}\right]-\mathbb{E}\left[\hat{\sigma}_{p}(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{\bm{Z}}))\;\middle|\;s,\bm{\kappa}_{s}\right]\right\|_{2}=o(1),$ (S34)
@@
$\left\|\mathrm{Cov}\left(\sigma(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{\bm{Z}}))\;\middle|\;s,\bm{\kappa}_{s}\right)-\mathrm{Cov}\left(\hat{\sigma}_{p}(\hat{{\bm{F}}}\text{vec}({\bm{H}}_{\bm{Z}}))\;\middle|\;s,\bm{\kappa}_{s}\right)\right\|_{2}=o(1),$ (S36)
@@
Finally, invoking the gradient decomposition and the corresponding bounds from Appendix C, together with the fact that ${\bm{P}}={\bm{S}}({\bm{S}}^{\top}{\bm{S}})^{-1}{\bm{S}}^{\top}$ has rank $O(1)$ and Assumptions 4.44.5, we conclude that both terms in (S39) vanish at the desired rate. In particular, the scaling in Assumption 4.5, the low-rank structure of $\bm{P}$}, and the spiked-covariance in Assumption 4.4 ensure that the contributions of the $\eta$-weighted gradient term and the spiked directions in $\mathrm{Cov}(\text{vec}({\bm{H}}_{\bm{Z}}))$ remain controlled, which completes the proof.